Project

General

Profile

NDOS Commissioning Meeting 2011/03/14

Attendance: Jon, Ralf, Bridgette, Peter, Andrew, Kurt, Mark Mathis, Tricia, Alena, Sue, Jarek, Greg, Rick K., Athans, Minerba, Enrique, Denis, Mat

Agenda:

  1. Installation/Channel Recovery Update
  2. PDB Work this past weekend
  3. DAQ Update & Plans
  4. DCS Status & Plans
  5. Shifters' report
  6. AOB

Installation/Channel Recovery:

  • secondary containment will be rolled out for scint. filling (for both truck and detector)
  • Minerba got list of people trained for filling, will organize filling shifts
  • 2 shifts/day (8am-3, 3-10pm ), 2 person crews, only one of which needs to be trained (oil, MAP, etc.)

Power Distribution System work (Ralf):

  • connected PDB on relay rack for muon catcher
  • PDB on side
  • ground strips on top of detector, but not on side
  • cable tray support for muon catcher need to be completed
  • will need a few more days to complete work

DAQ Hardware:

  • Andrew: new switch will be installed sometime this week, not clear yet which day
  • Leon: tool for noise-canceling boards was delivered at some front door at IU Cyclotron, but was lost at that point. Company sent more tools on Friday.
  • Mat: we now have 7 heat-sink assemblies with noise-canceling boards attached. We have 6 new APDs. Would like to install these 6 on the detector. There are some empty slots on top of diblock 2. Mat will organize the install for today.
  • Leon: more "normal" heatsinks will be sent this week to install cleaned APDs.
  • Neil and Rick K. have done some testing on fiber timing link. Hope to move to test stand this afternoon. Received some boards to evaluate new DCM hardware.
  • Status of spare DCMs: all seem to be working.
  • Peter: FEB checkout procedure? Sue will work on this.

DAQ Software:

  • Peter:
    • Friday evening we had a very hard time getting a stable run going. Looked like rate were continuing to drop over time. Suspects this was due to having power off on FEBs for an extended period.
    • We're ready to switch to using larger MTUs on DCMs and BufferNodes. Can be done via ifconfig, and will be done at boot/startup. These changes will be made today.
    • Other software updates will be pushed through today.
  • Kurt: still investigating lost trigger messages in DDS
  • Peter: investigating online monitoring code, will work with Alena on this.
  • Sue: made some performance upgrades that should improve user CPU usage. Will test on dcm-10 when it becomes available.

DCS: no update

Any other business: none